OpenStudy (anonymous):

Variable g is 8 more than variable w. Variable g is also 4 less than w. Which pair of equations best models the relationship between g and w? g = w − 8 g = w + 4 w = 8g w = g − 4 g = w + 8 g = w − 4 w = 8g w = g + 4

OpenStudy (anonymous):

g = w + 8 g = w − 4 Remember "is" means "equals," "more than" means you add, and "less than" mean you subtract.

OpenStudy (anonymous):

thxs can u help wit more

OpenStudy (anonymous):

means*

OpenStudy (anonymous):

What would most likely be the approximate height of the plant after 7 weeks? 4.8 centimeters 14.4 centimeters 16.5 centimeters 19.5 centimeters

OpenStudy (anonymous):

The graph is constant. You know what that means?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

As the number of weeks goes up, the height of the plant goes up as well.

OpenStudy (anonymous):

In week 5, what is the height of the plant?

OpenStudy (anonymous):

oh tru uhm... 4.8?

OpenStudy (anonymous):

As you notice the graph is gradually going up.

OpenStudy (anonymous):

In week 5, the height is 4.

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

If in week 5, the height is 4. Then in week 7, the height is closest to the height in week 5. What is the answer?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

16.5?

OpenStudy (anonymous):

Are the numbers 5 and 16.5 in the choices?

OpenStudy (anonymous):

Which is closest to 4?

OpenStudy (anonymous):

What would most likely be the approximate height of the plant after 7 weeks? 4.8 centimeters 14.4 centimeters 16.5 centimeters 19.5 centimeters

OpenStudy (anonymous):

4.8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

Which graph represents a positive linear association between x and y? Graph A Graph B Graph C Graph D